# Gradient Existence and Energy Finiteness of Local Minimizers in the Wasserstein $L^{\infty}$ Topology for Binary-Star Systems

### v1

@@ 4 Wasserstein $L^{\infty}$ Metric on ${R}\left(\mathbb{R}^{3}\right)$
Therefore, if $x\in\left\{x\mid\rho_{n}(x)\leq R\right\}\cap C\left(x_{n}\right)$, we get $\sigma_{n}(x)\leq\rho_{n}(x)+\frac{3}{4}\leq R+\frac{3}{4}<2R$, then we have $\sigma_{n}(x)\leq 2R$ for all $x\in\mathbb{R}^{3}$. One can also check $\int\rho_{n}=\int\sigma_{n}$ (notice $\left\{x\mid\rho_{n}(x)\geq 2R\right\}=\left\{x\mid\rho_{n}(x)\geq 2R\right\}\cap C\left(x_{n}\right)$). Since both $\rho_{n}$ and $\sigma_{n}$ are supported in $C\left(x_{n}\right)$ and $C\left(x_{n}\right)$ has diameter $\frac{\epsilon}{2}$, then diameter of $\operatorname{spt}\left(\rho_{n}-\sigma_{n}\right)$ is not larger than $\frac{\epsilon}{2}$. Thanks to Property ($i$), if $\int\rho_{n}=\int\sigma_{n}>0$, we have $W^{\infty}\left(\rho_{n},\sigma_{n}\right)\leq\frac{\epsilon}{2}$, where $W^{\infty}$ is the Wasserstein $L^{\infty}$ distance on $\mathcal{M}(\mathbb{R}^{3})$.
@@
Let $\sigma=\sum_{n=1}^{\infty}\sigma_{n}$, then $\sigma\in L^{\infty}\left(\mathbb{R}^{3}\right)$ and $\|\sigma\|_{L^{\infty}}\leq 2R$. We know $\int\sigma=\sum_{n}\int\sigma_{n}=\sum_{n}\int\rho_{n}={\int\rho}-{\int_{D}\rho}\stackrel{{\scriptstyle\mu\left(D\right)=0}}{{=}}\int\rho$. In particular, $\sigma$ has the same finite measure as $\rho$, and we can consider the Wasserstein $L^{\infty}$ metric between them.
Claim: $W^{\infty}(\rho,\sigma)<\epsilon$.
@@
In fact, suppose $\int\rho_{n}=\int\sigma_{n}>0$, since $W^{\infty}\left(\rho_{n},\sigma_{n}\right)\leq\frac{\epsilon}{2}$, we can set $\left({S,\Sigma,\nu}\right)=\left({C\left(x_{n}\right),M_{n},c_{n}\mu_{n}}\right)$, where $M_{n}$ is the collection of all Lebesgue measurable sets in $C\left(x_{n}\right)$, $c_{n}$ satisfies $c_{n}\mu_{n}\left({C\left(x_{n}\right)}\right)=c_{n}V=\int\rho_{n}=\int\sigma_{n}>0$, and $\mu_{n}$ satisfies $\mu_{n}(E)=\mu(E\cap C(x_{n}))$ for all measurable sets $E$ in $\mathbb{R}^{3}$, recall $\mu$ is the Lebesgue measure on $\mathbb{R}^{3}$. In particular, we know $c_{n}>0$, and $\mu_{n}(C(x_{n})^{\mathsf{c}})=0$, here $C(x_{n})^{\mathsf{c}}=\mathbb{R}^{3}\setminus C(x_{n})$ is the complement of $C(x_{n})$. Easy to verify that $\mu_{n}$ is a measure in $\mathbb{R}^{3}$ (see for example [17, Exercise 10 in Section 1.3]). Moreover, its restriction to $C(x_{n})$ coincides with the Lebesgue measure on $C(x_{n})$. Therefore, $\left({C\left(x_{n}\right),M_{n},c_{n}\mu_{n}}\right)$ is a (finite) measure space, with $c_{n}u_{n}\left({C\left(x_{n}\right)}\right)=\rho_{n}\left(\mathbb{R}^{3}\right)={\int\rho_{n}}$. Thanks to Remark 4.3, we can choose $\left({C\left(x_{n}\right),M_{n},c_{n}\mu_{n}}\right)$ as the measure space in the definition of $W^{\infty}\left(\rho_{n},\sigma_{n}\right)$ [^1] The reason for choosing such a measure space is to find transport maps $f_{n}$ and $g_{n}$ defined on $C(x_{n})$, which will facilitate the subsequent definition of $\widetilde{f_{n}}$, $\widetilde{g_{n}}$, $f$ and $g$, and the analysis of their properties., then we can find $f_{n}:C\left(x_{n}\right)\rightarrow\mathbb{R}^{3}$ and $g_{n}:C\left(x_{n}\right)\rightarrow\mathbb{R}^{3}$ with ${f_{n}}_{\#}(c_{n}\mu_{n})=\rho_{n},{g_{n}}_{\#}(c_{n}\mu_{n})=\sigma_{n}$, and
@@
Notice since $c_{n}>0$, we have $c_{n}\mu_{n}\ll\mu_{n},\mu_{n}\ll c_{n}\mu_{n}$, and then
@@
Therefore, given a measurable set $B\subset\mathbb{R}^{3}$, we have $c_{n}\mu_{n}\left({f_{n}^{-1}(B)}\right)=\rho_{n}(B)$. Since we know $\rho_{n}$ is supported in $C\left(x_{n}\right)$, then $\rho_{n}(B)=\rho_{n}\left({B\cap C\left(x_{n}\right)}\right)$. Therefore, $c_{n}\mu_{n}\left({f_{n}^{-1}(B)}\right)=\rho_{n}(B)=\rho_{n}\left({B\cap C\left(x_{n}\right)}\right)=c_{n}\mu_{n}\left({f_{n}^{-1}\left({B\cap\left({C\left(x_{n}\right)}\right)}\right)}\right)$. In particular, we know $c_{n}\mu_{n}\left({f_{n}^{-1}((C(x_{n}))^{\mathsf{c}})}\right)=c_{n}\mu_{n}\left({f_{n}^{-1}\left(\emptyset\right)}\right)=c_{n}\mu_{n}\left({\left(\emptyset\right)}\right)=0$, and then $\mu_{n}\left({f_{n}^{-1}((C(x_{n}))^{\mathsf{c}})}\right)=0$ since $c_{n}>0$. Then we can define $\widetilde{f_{n}}:C(x_{n})\rightarrow\mathbb{R}^{3}$ as the following:
@@
Notice by construction we know $0\in D$, so $0\notin C(x_{n})$ for any $n$. If $0\notin B$, we can check ${\widetilde{f_{n}}}^{-1}(B)=f_{n}^{-1}\left({B\cap C\left(x_{n}\right)}\right)$, then we have $c_{n}\mu_{n}\left({{\widetilde{f_{n}}}^{-1}(B)}\right)=c_{n}\mu_{n}\left({f_{n}^{-1}\left({B\cap C\left(x_{n}\right)}\right)}\right)=\rho_{n}(B)$. If $0\in B$, we can check ${\widetilde{f_{n}}}^{-1}(B)=f_{n}^{-1}\left({B\cap C\left(x_{n}\right)}\right)\cup f_{n}^{-1}\left({C\left(x_{n}\right)^{\mathsf{c}}}\right)$, and
@@
Take $\widetilde{U_{n}}\coloneq U_{n}\cup f_{n}^{-1}\left(\left({C\left(x_{n}\right)}\right)^{\mathsf{c}}\right)\cup g_{n}^{-1}\left(\left({C\left(x_{n}\right)}\right)^{\mathsf{c}}\right)$, since we know $\mu_{n}\left({f_{n}^{-1}((C(x_{n}))^{\mathsf{c}})}\right)=\mu_{n}\left({g_{n}^{-1}((C(x_{n}))^{\mathsf{c}})}\right)=0$, then $\mu_{n}\left(\widetilde{U_{n}}\right)=0$, and we have
@@
We define $\nu={\sum\limits_{n}{c_{n}\mu_{n}}}$, where $c_{n}$ satisfies $c_{n}\mu_{n}\left({C\left(x_{n}\right)}\right)=c_{n}V=\int\rho_{n}=\int\sigma_{n}$. Notice here we allow $c_{n}=0$. We can see $\nu$ is a measure on $\mathbb{R}^{3}$ and $\nu\left(\mathbb{R}^{3}\right)=\sum_{n}c_{n}\mu_{n}\left(\mathbb{R}^{3}\right)=\sum_{n}c_{n}\mu_{n}\left({C\left(x_{n}\right)}\right)=\sum_{n}\int\rho_{n}=\int\rho=\int\sigma=1<\infty$, thus $\nu$ is a finite measure.
@@
Then $\|f-g\|_{L^{\infty}\left(\mathbb{R}^{3}\right),\nu}\leq\sup\limits_{\{n|c_{n}>0\}}\left\|\widetilde{f_{n}}-\widetilde{g_{n}}\right\|_{L^{\infty}\left(C\left(x_{n}\right)\right),\mu_{n}}\leq\frac{3\epsilon}{4}$. Given a measurable set $B\subset\mathbb{R}^{3}$, let $B_{n}=B\cap C\left(x_{n}\right)$ be the measurable set in $C\left(x_{n}\right),B_{D}=B\cap D$, we know $\mu(D)=0$ implies $\mu\left(B_{D}\right)=0$. We also have if $n\in N_{0}$, then $\rho\left(B_{n}\right)=\rho_{n}\left(B_{n}\right)\leq\rho_{n}\left({C\left(x_{n}\right)}\right)={\int\rho_{n}}=0$, and then
$\rho(B)$ $=\int_{B}\rho=\int_{\cup_{n=1}^{\infty}B_{n}}\rho+\int_{B_{D}}\rho\stackrel{{\scriptstyle\mu\left(B_{D}\right)=0}}{{=}}\int_{\cup_{n=1}^{\infty}B_{n}}\rho$
$={\int_{\cup_{n\in N_{1}}B_{n}}\rho}={\sum_{n\in N_{1}}{\rho\left(B_{n}\right)}}={\sum_{n\in N_{1}}{\rho_{n}\left(B_{n}\right)}}$
$\stackrel{{\scriptstyle\widetilde{f_{n}}_{\#}(c_{n}\mu_{n})=\rho_{n}}}{{=}}{\sum_{n\in N_{1}}{c_{n}\mu_{n}\left({{\widetilde{f_{n}}}^{-1}\left(B_{n}\right)}\right)}}$
We can check $f^{-1}\left(\left\{0\right\}\right)={\bigcup\limits_{n\in N_{1}}{{\widetilde{f}}_{n}}^{-1}}\left(\left\{0\right\}\right)\cup\left\{0\right\}$. Then if $n\in N_{1}$, since ${\widetilde{f_{n}}}_{\#}\left({c_{n}\mu_{n}}\right)=\rho_{n}$, we have $c_{n}\mu_{n}\left({f^{-1}\left(\left\{0\right\}\right)}\right)=c_{n}\mu_{n}\left({{\widetilde{f_{n}}}^{-1}\left(\left\{0\right\}\right)}\right)=\rho_{n}\left(\left\{0\right\}\right)=0$. Easy to check if $n\in N_{0}$, $f^{-1}\left(\left\{0\right\}\right)\cap C\left(x_{n}\right)=\varnothing$, then $\mu_{n}\left({f^{-1}\left(\left\{0\right\}\right)}\right)=0$. Thus we know $v\left({f^{-1}\left\{0\right\}}\right)={\sum_{n}{c_{n}\mu_{n}\left({f^{-1}\left(\left\{0\right\}\right)}\right)}}=0$ (alternatively, this can also be seen by observing that $c_{n}=0$ for all $n\in N_{0}$).
@@
$\stackrel{{\scriptstyle c_{n}=0\text{ when }n\in N_{0}}}{{=}}{\sum_{n\in N_{1}}{c_{n}\mu_{n}\left({f^{-1}\left(B\right)}\right)}}+{\sum_{n\in N_{0}}{c_{n}\mu_{n}\left({f^{-1}\left(B\right)}\right)}}$
@@
Remark 4.8.
Lemma 4.7 (vi) facilitates showing that the local minimal energy $E_{J}(\rho)$ of the minimizer $\rho$ is always finite, and then taking derivatives makes sense. We will discuss more about it in Proposition 5.7.
In Theorem 2.25 we assume $(\rho,v)$ minimizes $E(\rho,v)$ locally on ${R}_{0}\left(\mathbb{R}^{3}\right)\times{V}\left(\mathbb{R}^{3}\right)$, which implies $\rho$ minimizes $E_{J}(\rho)$ locally on ${R}_{0}\left(\mathbb{R}^{3}\right)$ by Corollary 2.18. Some might doubt whether this condition is too weak to arrive at the conclusion Theorem 2.25 ($iv$) through (generalized) Lagrange multiplier theorem (Theorem A.4 and Remark A.5) if someone tries to follow the arguments of Auchmuty and Beals [5], where the centers of mass of admissible perturbations are not necessarily 0. But, in fact, it turns out that a local energy minimizer $\rho$ on ${R}_{0}\left(\mathbb{R}^{3}\right)$ is also stable under perturbations which shift its center of mass:
@@
There exists $\delta>0$ such that $E_{J}(\rho)\leq E_{J}(\kappa)$ whenever $\kappa\in{R}_{0}\left(\mathbb{R}^{3}\right)$ with $W^{\infty}(\rho,\kappa)<2\delta$. Now, suppose $\kappa\in{R}\left(\mathbb{R}^{3}\right)$ with $W^{\infty}(\rho,\kappa)<\frac{\delta}{2}$. Thanks to Lemma 4.7 $(ii)$, we know $spt~\kappa$ is also bounded. Lemma 4.7 $(iv)$ shows that $|x(\kappa)|<\frac{\delta}{2}$. Let $\widetilde{\kappa}(x)=\kappa(x+x(\kappa))$, then $\widetilde{\kappa}\in{R}_{0}\left(\mathbb{R}^{3}\right)$ and Lemma 4.7 $(iii)$ shows that $W^{\infty}(\widetilde{\kappa},\kappa)<\frac{\delta}{2}$. By triangle inequality (recall $W^{\infty}$ is a metric), we have $W^{\infty}(\widetilde{\kappa},\rho)\geq W^{\infty}(\widetilde{\kappa},\kappa)+W^{\infty}(\kappa,\rho)<\delta$, i.e. $\widetilde{\kappa}$ lies within $\delta$-neighborhood of $\rho$ in ${R}_{0}\left(\mathbb{R}^{3}\right)$. By translation invariance, $E_{J}(\rho)\leq E_{J}(\widetilde{\kappa})=E_{J}(\kappa)$. ∎
We should point out: suppose $\rho$ minimizes $E_{J}(\rho)$ locally on ${R}_{0}\left(\mathbb{R}^{3}\right)$ and let $\sigma\in L^{\infty}\left(\mathbb{R}^{3}\right)$. Even if the perturbation satisfies $\rho+t\sigma\in{R}\left(\mathbb{R}^{3}\right)$ for $t\in[0,1]$, it may not be $W^{\infty}$-continuous as function of $t$, and we do not know if $E_{J}(\rho+t\sigma)\geq E_{J}(\rho)$ for $t\in[0,1]$; nevertheless, when $\sigma$ is supported on a set which diameter is small enough, with $\int\sigma=0$, then $W^{\infty}(\rho,\rho+t\sigma)$ is small enough for all $t$ thanks to Lemma 4.7 $(i)$, hence $\rho$ minimizes $E_{J}(\rho+t\sigma)$ for all $t$. Therefore, $\sigma$ will then be a useful variation of $E_{J}(\rho)$.

### v2

@@ 4 Wasserstein $L^{\infty}$ Metric on ${R}\left(\mathbb{R}^{3}\right)$
Then we have $\sigma_{n}(x)\leq 2R$ for all $x\in\mathbb{R}^{3}$. One can also check $\int_{\mathbb{R}^{3}}\rho_{n}\,dx=\int_{\mathbb{R}^{3}}\sigma_{n}\,dx$ (notice $\left\{x\mid\rho_{n}(x)\geq 2R\right\}=\left\{x\mid\rho_{n}(x)\geq 2R\right\}\cap C\left(x_{n}\right)$). Since both $\rho_{n}$ and $\sigma_{n}$ are supported in $C\left(x_{n}\right)$ and $C\left(x_{n}\right)$ has diameter $\frac{\epsilon}{2}$, then diameter of $\operatorname{spt}\left(\rho_{n}-\sigma_{n}\right)$ is not larger than $\frac{\epsilon}{2}$. Thanks to Property ($i$), if $\int_{\mathbb{R}^{3}}\rho_{n}\,dx=\int_{\mathbb{R}^{3}}\sigma_{n}\,dx>0$, we have $W^{\infty}\left(\rho_{n},\sigma_{n}\right)\leq\frac{\epsilon}{2}$, where $W^{\infty}$ is the Wasserstein $L^{\infty}$ distance on $\mathcal{M}(\mathbb{R}^{3})$.
Let $\sigma=\sum_{n=1}^{\infty}\sigma_{n}$, then $\sigma\in L^{\infty}\left(\mathbb{R}^{3}\right)$ and $\|\sigma\|_{L^{\infty}}\leq 2R$. We know
$\int_{\mathbb{R}^{3}}\sigma\,dx=\sum_{n}\int_{\mathbb{R}^{3}}\sigma_{n}\,dx=\sum_{n}\int_{\mathbb{R}^{3}}\rho_{n}\,dx={\int_{\mathbb{R}^{3}}\rho\,dx}-{\int_{D}\rho\,dx}\stackrel{{\scriptstyle\mu\left(D\right)=0}}{{=}}\int_{\mathbb{R}^{3}}\rho\,dx$
@@
In particular, $\sigma$ has the same finite measure as $\rho$, and we can now estimate the Wasserstein $L^{\infty}$ metric between them. We claim: $W^{\infty}(\rho,\sigma)<\epsilon$.
In order to show the claim $W^{\infty}(\rho,\sigma)<\epsilon$, we first consider if $\int_{\mathbb{R}^{3}}\rho_{n}\,dx=\int_{\mathbb{R}^{3}}\sigma_{n}\,dx>0$, since $W^{\infty}\left(\rho_{n},\sigma_{n}\right)\leq\frac{\epsilon}{2}$, we can set $\left({S,\Sigma,\nu}\right)=\left({C\left(x_{n}\right),M_{n},c_{n}\mu_{n}}\right)$, where $M_{n}$ is the collection of all Lebesgue measurable sets in $C\left(x_{n}\right)$, $c_{n}$ satisfies
$c_{n}\mu_{n}\left({C\left(x_{n}\right)}\right)=c_{n}V=\int_{\mathbb{R}^{3}}\rho_{n}\,dx=\int_{\mathbb{R}^{3}}\sigma_{n}\,dx>0$
@@
and $\mu_{n}$ satisfies $\mu_{n}(E)=\mu(E\cap C(x_{n}))$ for all measurable sets $E$ in $\mathbb{R}^{3}$, recall $\mu$ is the Lebesgue measure on $\mathbb{R}^{3}$. In particular, we know $c_{n}>0$, and $\mu_{n}(C(x_{n})^{\mathsf{c}})=0$, here $C(x_{n})^{\mathsf{c}}=\mathbb{R}^{3}\setminus C(x_{n})$ is the complement of $C(x_{n})$. Easy to verify that $\mu_{n}$ is a measure in $\mathbb{R}^{3}$ (see for example [17, Exercise 10 in Section 1.3]). Moreover, its restriction to $C(x_{n})$ coincides with the Lebesgue measure on $C(x_{n})$. Therefore, $\left({C\left(x_{n}\right),M_{n},c_{n}\mu_{n}}\right)$ is a (finite) measure space, with $c_{n}u_{n}\left({C\left(x_{n}\right)}\right)=\rho_{n}\left(\mathbb{R}^{3}\right)={\int_{\mathbb{R}^{3}}\rho_{n}\,dx}$. Thanks to Remark 4.3, we can choose $\left({C\left(x_{n}\right),M_{n},c_{n}\mu_{n}}\right)$ as the measure space in the definition of $W^{\infty}\left(\rho_{n},\sigma_{n}\right)$ [^1] The reason for choosing such a measure space is to find transport maps $f_{n}$ and $g_{n}$ defined on $C(x_{n})$, which will facilitate the subsequent definition of $\widetilde{f_{n}}$, $\widetilde{g_{n}}$, $f$ and $g$, and the analysis of their properties., then we can find $f_{n}:C\left(x_{n}\right)\rightarrow\mathbb{R}^{3}$ and $g_{n}:C\left(x_{n}\right)\rightarrow\mathbb{R}^{3}$ with ${f_{n}}_{\#}(c_{n}\mu_{n})=\rho_{n},{g_{n}}_{\#}(c_{n}\mu_{n})=\sigma_{n}$, and
@@
Notice since $c_{n}>0$, we have $c_{n}\mu_{n}\ll\mu_{n}$ and $\mu_{n}\ll c_{n}\mu_{n}$, and then
@@
Therefore, given a measurable set $B\subset\mathbb{R}^{3}$, we have $c_{n}\mu_{n}\left({f_{n}^{-1}(B)}\right)=\rho_{n}(B)$. Since we know $\rho_{n}$ is supported in $C\left(x_{n}\right)$, then $\rho_{n}(B)=\rho_{n}\left({B\cap C\left(x_{n}\right)}\right)$. Therefore,
$c_{n}\mu_{n}\left({f_{n}^{-1}(B)}\right)=\rho_{n}(B)=\rho_{n}\left({B\cap C\left(x_{n}\right)}\right)=c_{n}\mu_{n}\left({f_{n}^{-1}\left({B\cap\left({C\left(x_{n}\right)}\right)}\right)}\right)$
In particular, we know
$c_{n}\mu_{n}\left({f_{n}^{-1}((C(x_{n}))^{\mathsf{c}})}\right)=c_{n}\mu_{n}\left({f_{n}^{-1}\left(\emptyset\right)}\right)=c_{n}\mu_{n}\left({\left(\emptyset\right)}\right)=0$
and then $\mu_{n}\left({f_{n}^{-1}((C(x_{n}))^{\mathsf{c}})}\right)=0$ since $c_{n}>0$. Then we can define $\widetilde{f_{n}}:C(x_{n})\rightarrow\mathbb{R}^{3}$ as the following:
@@
Notice by construction we know $0\in D$, so $0\notin C(x_{n})$ for any $n$. If $0\notin B$, we can check ${\widetilde{f_{n}}}^{-1}(B)=f_{n}^{-1}\left({B\cap C\left(x_{n}\right)}\right)$, then we have
$c_{n}\mu_{n}\left({{\widetilde{f_{n}}}^{-1}(B)}\right)=c_{n}\mu_{n}\left({f_{n}^{-1}\left({B\cap C\left(x_{n}\right)}\right)}\right)=\rho_{n}(B)$
If $0\in B$, we can check ${\widetilde{f_{n}}}^{-1}(B)=f_{n}^{-1}\left({B\cap C\left(x_{n}\right)}\right)\cup f_{n}^{-1}\left({C\left(x_{n}\right)^{\mathsf{c}}}\right)$, and
@@
Take $\widetilde{U_{n}}\coloneq U_{n}\cup f_{n}^{-1}\left(\left({C\left(x_{n}\right)}\right)^{\mathsf{c}}\right)\cup g_{n}^{-1}\left(\left({C\left(x_{n}\right)}\right)^{\mathsf{c}}\right)$, since we know
$\mu_{n}\left({f_{n}^{-1}((C(x_{n}))^{\mathsf{c}})}\right)=\mu_{n}\left({g_{n}^{-1}((C(x_{n}))^{\mathsf{c}})}\right)=0$
then $\mu_{n}\left(\widetilde{U_{n}}\right)=0$, and we have
@@
We define $\nu={\sum\limits_{n}{c_{n}\mu_{n}}}$, where $c_{n}$ satisfies $c_{n}\mu_{n}\left({C\left(x_{n}\right)}\right)=c_{n}V=\int_{\mathbb{R}^{3}}\rho_{n}\,dx=\int_{\mathbb{R}^{3}}\sigma_{n}\,dx$. Notice here we allow $c_{n}=0$. We can see $\nu$ is a measure on $\mathbb{R}^{3}$ and
$\nu\left(\mathbb{R}^{3}\right)=\sum_{n}c_{n}\mu_{n}\left(\mathbb{R}^{3}\right)=\sum_{n}c_{n}\mu_{n}\left({C\left(x_{n}\right)}\right)=\sum_{n}\int_{\mathbb{R}^{3}}\rho_{n}\,dx=\int_{\mathbb{R}^{3}}\rho\,dx=\int_{\mathbb{R}^{3}}\sigma\,dx=1<\infty$
Thus $\nu$ is a finite measure.
@@
Then
$\|f-g\|_{L^{\infty}\left(\mathbb{R}^{3}\right),\nu}\leq\sup\limits_{\{n|c_{n}>0\}}\left\|\widetilde{f_{n}}-\widetilde{g_{n}}\right\|_{L^{\infty}\left(C\left(x_{n}\right)\right),\mu_{n}}\leq\frac{3\epsilon}{4}$
Given a measurable set $B\subset\mathbb{R}^{3}$, let $B_{n}=B\cap C\left(x_{n}\right)$ be the measurable set in $C\left(x_{n}\right),B_{D}=B\cap D$, we know $\mu(D)=0$ implies $\mu\left(B_{D}\right)=0$. We also have if $n\in N_{0}$, then
$\rho\left(B_{n}\right)=\rho_{n}\left(B_{n}\right)\leq\rho_{n}\left({C\left(x_{n}\right)}\right)={\int_{\mathbb{R}^{3}}\rho_{n}\,dx}=0$
and then
$\rho(B)$ $=\int_{B}\rho\,dx$
$=\int_{\cup_{n=1}^{\infty}B_{n}}\rho\,dx+\int_{B_{D}}\rho\,dx$
$=\int_{\cup_{n=1}^{\infty}B_{n}}\rho\,dx\qquad(\text{since }\mu(B_{D})=0)$
$=\int_{\cup_{n\in N_{1}}B_{n}}\rho\,dx$
$=\sum_{n\in N_{1}}\rho(B_{n})$
$=\sum_{n\in N_{1}}\rho_{n}(B_{n})$
$=\sum_{n\in N_{1}}c_{n}\mu_{n}\bigl(\widetilde{f_{n}}^{-1}(B_{n})\bigr)\qquad(\text{since }\widetilde{f_{n}}_{\#}(c_{n}\mu_{n})=\rho_{n})$
We can check $f^{-1}\left(\left\{0\right\}\right)={\bigcup\limits_{n\in N_{1}}{{\widetilde{f}}_{n}}^{-1}}\left(\left\{0\right\}\right)\cup\left\{0\right\}$. Then if $n\in N_{1}$, since ${\widetilde{f_{n}}}_{\#}\left({c_{n}\mu_{n}}\right)=\rho_{n}$, we have
$c_{n}\mu_{n}\left({f^{-1}\left(\left\{0\right\}\right)}\right)=c_{n}\mu_{n}\left({{\widetilde{f_{n}}}^{-1}\left(\left\{0\right\}\right)}\right)=\rho_{n}\left(\left\{0\right\}\right)=0$
Easy to check if $n\in N_{0}$, $f^{-1}\left(\left\{0\right\}\right)\cap C\left(x_{n}\right)=\varnothing$, then $\mu_{n}\left({f^{-1}\left(\left\{0\right\}\right)}\right)=0$. Thus we know $v\left({f^{-1}\left\{0\right\}}\right)={\sum_{n}{c_{n}\mu_{n}\left({f^{-1}\left(\left\{0\right\}\right)}\right)}}=0$ (alternatively, this can also be seen by observing that $c_{n}=0$ for all $n\in N_{0}$).
@@
$={\sum_{n\in N_{1}}{c_{n}\mu_{n}\left({f^{-1}\left(B\right)}\right)}}+{\sum_{n\in N_{0}}{c_{n}\mu_{n}\left({f^{-1}\left(B\right)}\right)}}\qquad(\text{since }c_{n}=0\text{ when }n\in N_{0})$
@@
In Theorem 2.22 we assume $(\rho,v)$ minimizes $E(\rho,v)$ locally on ${R}_{0}\left(\mathbb{R}^{3}\right)\times{V}\left(\mathbb{R}^{3}\right)$, which implies $\rho$ minimizes $E_{J}(\rho)$ locally on ${R}_{0}\left(\mathbb{R}^{3}\right)$ by Corollary 2.18. Although in the paper of Auchmuty and Beals [5], the center of mass after perturbation is not necessarily zero, it turns out that a local energy minimizer $\rho$ on ${R}_{0}\left(\mathbb{R}^{3}\right)$ is also stable under perturbations which shift its center of mass:
@@
There exists $\delta>0$ such that $E_{J}(\rho)\leq E_{J}(\kappa)$ whenever $\kappa\in{R}_{0}\left(\mathbb{R}^{3}\right)$ with $W^{\infty}(\rho,\kappa)<2\delta$. Now, suppose $\kappa\in{R}\left(\mathbb{R}^{3}\right)$ with $W^{\infty}(\rho,\kappa)<\frac{\delta}{2}$. Thanks to Lemma 4.5 $(ii)$, we know $spt~\kappa$ is also bounded. Lemma 4.5 $(iv)$ shows that $|x(\kappa)|<\frac{\delta}{2}$. Let $\widetilde{\kappa}(x)=\kappa(x+x(\kappa))$, then $\widetilde{\kappa}\in{R}_{0}\left(\mathbb{R}^{3}\right)$ and Lemma 4.5 $(iii)$ shows that $W^{\infty}(\widetilde{\kappa},\kappa)<\frac{\delta}{2}$. By triangle inequality (recall $W^{\infty}$ is a metric), we have $W^{\infty}(\widetilde{\kappa},\rho)\leq W^{\infty}(\widetilde{\kappa},\kappa)+W^{\infty}(\kappa,\rho)<\delta$, i.e. $\widetilde{\kappa}$ lies within $\delta$-neighborhood of $\rho$ in ${R}_{0}\left(\mathbb{R}^{3}\right)$. By translation invariance, $E_{J}(\rho)\leq E_{J}(\widetilde{\kappa})=E_{J}(\kappa)$. ∎
We should point out: suppose $\rho$ minimizes $E_{J}(\rho)$ locally on ${R}_{0}\left(\mathbb{R}^{3}\right)$ and let $\sigma\in L^{\infty}\left(\mathbb{R}^{3}\right)$. Even if the perturbation satisfies $\rho+t\sigma\in{R}\left(\mathbb{R}^{3}\right)$ for $t\in[0,1]$, it may not be $W^{\infty}$-continuous as function of $t$, and we do not know if $E_{J}(\rho+t\sigma)\geq E_{J}(\rho)$ for $t\in[0,1]$; nevertheless, when $\sigma$ is supported on a set which diameter is small enough, with $\int_{\mathbb{R}^{3}}\sigma\,dx=0$, then $W^{\infty}(\rho,\rho+t\sigma)$ is small enough for all $t$ thanks to Lemma 4.5 $(i)$, hence $\rho$ minimizes $E_{J}(\rho+t\sigma)$ for all $t$. Therefore, $\sigma$ will then be a useful variation of $E_{J}(\rho)$.